OpenStudy (anonymous):

A solution contains 11 milliliters of alcohol and 14 milliliters of water. If another solution is to have the same concentration of alcohol in water but is to contain 28 milliliters of water, how much alcohol must it contain?

OpenStudy (anonymous):

11/14 = x/28 solve for x

OpenStudy (anonymous):

@Hero , I multiply both sides by 28, right?

hero (hero):

And remember than 28/14 = 2

hero (hero):

that*

OpenStudy (anonymous):

and I multiply that by the 11?

hero (hero):

Correct

OpenStudy (anonymous):

22!

OpenStudy (anonymous):

Thank you!